# Thermal and non-thermal dark matters with gravitational neutrino reheating

### v1

@@ Appendix A Construction of $y_{ij}$ using Casas-Ibarra Parametrization
where $z=a+ib$ is a complex angle. The diagonal light neutrino mass matrix $m^{\rm d}_{\rm\nu}$ is calculable using the best-fit values of solar and atmospheric mass obtained from the latest neutrino oscillation data Esteban:2018azc . The elements of Yukawa coupling matrix $y$ for a specific value of $z$ can be obtained for different choices of the heavy neutrino masses. For example, with $M_{1}=3\times 10^{12}$ GeV,$M_{2}=5\times 10^{13}$ GeV, $\{a,b\}=\{2.5,1.4\}$,$m_{3}=0.05$ eV,$m_{2}=0.008$ eV and taking $M_{3}$, $m_{1}$ are the free parameter, we obtain
@@
Appendix B Listed the DM annihilation cross-section

### v2

@@ Appendix A Construction of $y_{ij}$ using Casas-Ibarra Parametrization
where $z=a+ib$ is a complex angle. The diagonal light neutrino mass matrix $m^{\rm d}_{\rm\nu}$ is calculable using the best-fit values of solar and atmospheric mass obtained from the latest neutrino oscillation data Esteban:2018azc . The elements of Yukawa coupling matrix $y$ for a specific value of $z$ can be obtained for different choices of the heavy neutrino masses. For example, with $M_{1}=3\times 10^{12}$ GeV,$M_{2}=5\times 10^{13}$ GeV, $\{a,b\}=\{2.5,1.4\}$, using normal hierarchy, $m_{3}=0.05$ eV,$m_{2}=0.008$ eV and taking $M_{3}$, $m_{1}$ are the free parameter, we obtain
@@
Similarly, we can construct the Yukawa coupling $y$ using the inverted hierarchy (IH). The primary distinction lies in the choice of the orthogonal matrix, where the form of orthogonal matrix $\mathbb{R}$ is,
[…]
For IH, the neutrino masses are given by $m_{2}=0.05$ eV, $m_{1}=0.0497$ eV, and $m_{3}$ is treated as a free parameter. Using these values, we obtain the following Yukawa coupling,
$y=\begin{pmatrix}-0.044-0.112\,i&\quad-0.402+0.125\,i&\quad(0.111-0.100\,i)\,\sqrt{m_{3}\,M_{3}}/v\\ 0.073-0.032\,i&\quad-0.120-0.199\,i&\quad\quad 0.738\,\sqrt{m_{3}\,M_{3}}/v\\ -0.107+0.024\,i&\quad 0.086+0.300\,i&\quad 0.657\,\sqrt{m_{3}\,M_{3}}/v\end{pmatrix}.$ (91)
Therefore, for IH, $y_{\rm i3}\propto\sqrt{m_{3}\,M_{3}}/v$ i.e, $\beta^{2}\propto m_{3}.$
Appendix B RG Effects in Neutrino Yukawa coupling
The one-loop renormalization group evolution (RGE) equations of the relevant parameters associated with the Seesaw-I model are :
The gauge couplings:
[…]
The Higgs quadratic and quartic couplings:
$16\,\pi^{2}\frac{d\mu^{2}_{H}}{d\ln\mu}=\left(-\frac{9}{10}g_{1}^{2}-\frac{9}{2}g_{2}^{2}+12\lambda+2\,T\right)\,\mu^{2}_{H}$
$116\,\pi^{2}\frac{d\lambda_{H}}{d\ln\mu}=\left(-\frac{9}{5}g_{1}^{2}-9g_{2}^{2}+24\lambda_{H}+4T\right)\,\lambda_{H}+\frac{27}{200}g_{1}^{4}+\frac{9}{20}g_{1}^{2}g_{2}^{2}+\frac{9}{8}g_{2}^{4}-2T$ 1
The Yukawa couplings:
$16\,\pi^{2}\frac{dY_{u}}{d\ln\mu}=\left(-\frac{17}{20}g_{1}^{2}-\frac{9}{4}g_{2}^{2}-8g_{3}^{2}+T\right)Y_{u}+\frac{3}{2}(Y_{u}Y_{u}^{\dagger}Y_{u})-\frac{3}{2}(Y_{u}Y_{d}^{\dagger}Y_{d})$
$16\,\pi^{2}\frac{dY_{d}}{d\ln\mu}=-\frac{1}{4}\left(g_{1}^{2}+9g_{2}^{2}+32g_{3}^{2}-4\,T\right)Y_{d}+\frac{3}{2}(Y_{d}Y_{d}^{\dagger}Y_{d})-\frac{3}{2}(Y_{d}Y_{u}^{\dagger}Y_{u})$
$16\,\pi^{2}\frac{dY_{e}}{d\ln\mu}=\left(-\frac{9}{4}g_{1}^{2}-\frac{9}{4}g_{2}^{2}+T\right)Y_{e}+\frac{3}{2}(Y_{e}Y_{e}^{\dagger}Y_{e})-\frac{3}{2}(Y_{e}Y_{\nu}^{\dagger}Y_{\nu})$
$16\,\pi^{2}\frac{dY_{\nu}}{d\ln\mu}=\left(-\frac{9}{20}g_{1}^{2}-\frac{9}{4}g_{2}^{2}+T\right)Y_{\nu}+\frac{3}{2}(Y_{\nu}Y_{\nu}^{\dagger}Y_{\nu})-\frac{3}{2}(Y_{\nu}Y_{e}^{\dagger}Y_{e})$
The Majorana mass:
$16\,\pi^{2}\frac{dM_{\nu}}{d\ln\mu}=2M_{v}(Y_{v}Y_{v}^{\dagger})^{T}+Y_{v}Y_{v}^{\dagger}M_{v}$
where $T=\text{Tr}\left(3Y_{d}Y_{d}^{\dagger}+3Y_{u}Y_{u}^{\dagger}+Y_{e}Y_{e}^{\dagger}+Y_{\nu}Y_{\nu}^{\dagger}\right)$ and $T1=\text{Tr}\left(3(Y_{d}Y_{d}^{\dagger})^{2}+3(Y_{u}Y_{u}^{\dagger})^{2}+(Y_{e}Y_{e}^{\dagger})^{2}+(Y_{\nu}Y_{\nu}^{\dagger})^{2}\right)$ SARAH provides with a numerical solution to the derived renormalization group equations. In the figure 13, we have shown ratio of running of $|y_{33}|$ and $|y_{22}|$ with the energy scale $\mu$. Initially, $\frac{|y_{33}|}{|y_{22}|}=9.4\times 10^{-10}$, after taking running, $\frac{|y_{33}|}{|y_{22}|}\simeq 9.9\times 10^{-10}$ at $\mu=4$ MeV.
Figure 13: Evolution of $\frac{|y_{33}|}{|y_{22}|}$ with energy scale $\mu=T^{\rm max}_{\rm rad}$ to $\mu=T_{\rm BBN}=4$ MeV.
Appendix C Listed the DM annihilation cross-section